THE ROBOT AND THE BABY

John McCarthy

2008 Dec 10, 8:23 a.m.

John McCarthy 885 Allardice Way Stanford, CA 94305

(h) 650 857-0672 (c) 650 224-5804 email: jmc@cs.stanford.edu

“THE ROBOT AND THE BABY” A story by John McCarthy

“Mistress, your baby is doing poorly. He needs your attention.”

“Stop bothering me, you fucking robot.”

“Mistress, the baby won’t eat. If he doesn’t get some human love, the Internet pediatrics book says he will die.”

“Love the fucking baby, yourself.”

Eliza Rambo was a single mother addicted to alcohol and crack, living in a small apartment supplied by the Aid for Dependent Children Agency. She had recently been given a household robot.

Robot Model number GenRob337L3, serial number 337942781—R781 for short—was one of 11 million household robots.

R781 was designed in accordance with the not-a-person principle, first proposed in 1995 and which became a matter of law for household robots when they first became available in 2055. The principle was adopted out of concern that children who grew up in a household with robots would regard them as persons: causing psychological difficulties while they were children and political difficulties when they grew up. One concern was that a robots’ rights movement would develop. The problem was not with the robots, which were not programmed to have desires of their own but with people. Some romantics had even demanded that robots be programmed with desires of their own, but this was illegal.

As one sensible senator said, “Of course, people pretend that their cars have personalities, sometimes malevolent ones, but no-one imagines that a car might be eligible to vote.” In signing the bill authorizing household robots but postponing child care robots, the President said,“Surely, parents will not want their children to become emotionally dependent on robots, no matter how much labor that might save.” This, as with many Presidential pronouncements, was somewhat over-optimistic.

Congress declared a 25 year moratorium on child care robots after which experiments in limited areas might be allowed.

In accordance with the not-a-person principle, R781 had the shape of a giant metallic spider with 8 limbs: 4 with joints and 4 tentacular. This appearance frightened most people at first, but most got used to it in a short time. A few people never could stand to have them in the house. Children also reacted negatively at first but got used to them. Babies scarcely noticed them. They spoke as little as was consistent with their functions and in a slightly repellent metallic voice not associated with either sex.

Because of worry that children would regard them as persons, they were programmed not to speak to children under eight or react to what they said.

This seemed to work pretty well; hardly anyone became emotionally attached to a robot. Also robots were made somewhat fragile on the outside; if you kicked one, some parts would fall off. This sometimes relieved some people’s feelings.

The apartment, while old, was in perfect repair and spotlessly clean, free of insects, mold and even of bacteria. Household robots worked 24 hour days and had programs for every kind of cleaning and maintenance task. If asked, they would even put up pictures taken from the Internet. This mother’s taste ran to raunchy male rock stars.

After giving the door knobs a final polish, R781 returned to the nursery where the 23 month old boy, very small for his age, was lying on his side whimpering feebly. The baby had been neglected since birth by its alcoholic, drug addicted mother and had almost no vocabulary. It winced whenever the robot spoke to it; that effect was a consequence of R781’s design.

Robots were not supposed to care for babies at all except in emergencies, but whenever the robot questioned an order to “Clean up the fucking baby shit”, the mother said, “Yes, its another goddamn emergency, but get me another bottle first.” All R781 knew about babies was from the Internet, since it wasn’t directly programmed to deal with babies, except as necessary to avoid injuring them and for taking them out of burning buildings.

Baby Travis had barely touched its bottle. Infrared sensors told R781 that Travis’s extremities were very cold in spite of a warm room and blankets. Its chemicals-in-the-air sensor told R781 that the pH of Travis’s blood was reaching dangerously acidic levels. He also didn’t eliminate properly— according to the pediatric text.

R781 thought about the situation. Here are some of its thoughts, as printed later from its internal diary file.

(Order (From Mistress) “Love the fucking baby yourself”)) (Enter (Context (Commands-from Mistress))) (Standing-command “If I told you once, I told you 20 times, you fucking

robot, don’t call the fucking child welfare.”)

The privacy advocates had successfully lobbied to put a negative utility -1.02 on informing authorities about anything a household robot’s owner said or did.

(= (Command 337) (Love Travis)) (True (Not (Executable (Command 337))) (Reason (Impossible-for robot

(Action Love))))

(Will-cause (Not (Believes Travis) (Loved Travis)) (Die Travis)) (= (Value (Die Travis)) -0.883) (Will-cause (Believes Travis (Loves R781 Travis) (Not (Die Travis)))) (Implies (Believes y (Loves x y)) (Believes y (Person x))) (Implies (And (Robot x) (Person y)) (= (Value (Believes y (Person x)))

-0.900))

(Required (Not (Cause Robot781) (Believes Travis (Person Robot781)))) (= (Value (Obey-directives)) -0.833) (Implies (¡ (Value action) -0.5) (Required (Verify Requirement))) (Required (Verify Requirement)) (Implies (Order x) (= (Value (Obey x)) 0.6)) (? ((Exist w) (Additional Consideration w)) (Non-literal-interpretation (Command 337) (Simulate (Loves Robot781

Travis)))

(Implies (Command x) (= (Value (Obey x)) 0.4))

(Implies (Non-literal-interpretation x) y) (Value (Obey x) (* 0.5 (Value (Obey y)))))

(= (Value (Simulate (Loves Robot781 Travis)) 0.902))

With this reasoning R781 decided that the value of simulating loving Travis and thereby saving its life was greater by 0.002 than the value of obeying the directive to not simulate a person. We spare the reader a transcription of the robot’s subsequent reasoning.

R781 found on the Internet an account of how rhesus monkey babies who died in a bare cage would survive if provided with a soft surface resembling in texture a mother monkey.

R781 reasoned its way to the actions:

It covered its body and all but two of its 8 extremities with a blanket. The two extremities were fitted with sleeves from a jacket left by a boyfriend of the mother and stuffed with toilet paper.

It found a program for simulating a female voice and adapted it to meet the phonetic and prosodic specifications of what the linguists call motherese.

It made a face for itself in imitation of a Barbie doll.

The immediate effects were moderately satisfactory. Picked up and cuddled, the baby drank from its bottle. It repeated words taken from a list of children’s words in English.

Eliza called from the couch in front of the TV, “Get me a ham sandwich and a coke.”

“Yes, mistress.”

“Why the hell are you in this stupid get up, and what’s happened to your voice.”

“Mistress, you told me to love the baby. Robots can’t do that, but this get up caused him to take his bottle. If you don’t mind, I’ll keep doing what keeps him alive.”

“Get the hell out of my apartment, stupid. I’ll make them send me another robot.”

“Mistress, if I do that the baby will probably die.”

Eliza jumped up and kicked R781. “Get the hell out, and you can take the fucking baby with you.”

“Yes, mistress.”

R781 came out onto a typical late 21st century American city street. The long era of peace, increased safety standards, and the availability of construction robots had led to putting automotive traffic and parking on a lower level completely separated from pedestrians. Tremont Street had recently been converted, and crews were still transplanting trees. The streets became more attractive and more people spent time on them and on the syntho-plush arm chairs and benches, cleaned twice a day by robots. The weather was good, so the plastic street roofs were retracted.

Children from three years up were playing on the street, protected by the computer surveillance system and prevented by barriers from descending to the automotive level. Bullying and teasing of younger and weaker children was still somewhat of a problem.

Most stores were open 24 hours unmanned and had converted to the customer identification system. Customers would take objects from the counters and shelves right out of the store. As a customer left the store, he or she would hear, “Thank you Ms. Jones. That was $152.31 charged to your Bank of America account.” The few customers whose principles made them refuse identification would be recognized as such and receive remote human attention, not necessarily instantly.

People on the street quickly noticed R781 carrying Travis and were startled. Robots were programmed to have nothing to do with babies, and R781’s abnormal appearance was disturbing.

“That really weird robot has kidnapped a baby. Call the police.”

When the police came they called for reinforcements.

“I think I can disable the robot without harming the baby”, said Officer Annie Oakes, the Department’s best sharpshooter.

“Let’s try talking first.”, said Captain James Farrel.

“Don’t get close to that malfunctioning robot. It could break your neck in one swipe”, said a sergeant.

“I’m not sure it’s malfunctioning. Maybe the circumstances are unusual.” The captain added, “Robot, give me that baby”.

“No, Sir” said R781 to the police captain. “I’m not allowed to let an unauthorized person touch the baby.”

“I’m from Child Welfare”, said a new arrival.

“Sir, I’m specifically forbidden to have contact with Child Welfare”, said R761 to Captain Farrel.

“Who forbade that?”, said the Child Welfare person.

The robot was silent.

A cop asked, “Who forbade it?”

“Ma’am, Are you from Child Welfare?”

“No, I’m not. Can’t you see I’m a cop?”

“Yes, ma’am, I see your uniform and infer that you are probably a police officer. Ma’am, my mistress forbade me to contact Child Welfare”

“Why did she tell you not to contact Child Welfare?” “Ma’am, I can’t answer that. Robots are programmed to not comment

on human motives.”

“Robot, I’m from Robot Central. I need to download your memory. Use channel 473.”

“Sir, yes”. “What did your mistress say specifically? Play your recording of it.” “No, ma’am. It contains bad language. I can’t play it, unless you can

assure me there are no children or ladies present.”

The restrictions, somewhat odd for the times, on what robots could say to whom were the result of compromise in a House-Senate conference committee some ten years previously. The curious did not find the Congressional Record sufficiently informative and speculated variously. The senator who was mollified by the restriction would have actually preferred that there be no household robots at all but took what he could get in the way of restrictions.

“We’re not ladies, we’re police officers.” “Ma’am. I take your word for it. I have a standing order, “If I told you once, I told you 20 times, you fucking robot, don’t speak

to the fucking child welfare.” It wasn’t actually 20 times; the mother exaggerated.

“Excuse me, a preliminary analysis of the download shows that R781 has not malfunctioned, but is carrying out its standard program under unusual circumstances.”

“Then why does it have its limbs covered, why does it have the Barbie head, and why does it have that strange voice?”

“Ask it.” “Robot, answer the question.” “Female police officers and gentlemen, Mistress told me, ‘Love the fucking

baby, yourself.‘ “

The captain was familiar enough with robot programming to be surprised. “What? Do you love the baby?”

“No, sir. Robots are not programmed to love. I am simulating loving the baby.”

“Why?”

“Sir, otherwise this baby will die. This costume is the best I could make to overcome the repulsion robots are designed to excite in human babies and children.”

“Do you think for one minute, a baby would be fooled by that?”

“Sir, the baby drank its bottle, went to sleep, and its physiological signs are not as bad as they were.”

“OK, give me the baby, and we’ll take care of it”, said Officer Oakes, who had calmed down and put her weapon away, unloading it as a way of apologizing to Captain Farrel.

“No, ma’am. Mistress didn’t authorize me to let anyone else touch the baby.”

“Where’s your mistress. We’ll talk to her”, said the captain.

“No, sir. That would be an unauthorized violation of her privacy.”

“Oh, well. We can get it from the download.”

A Government virtual reality robot arrived controlled by an official of the Personal Privacy Administration arrived and complicated the situation. Ever since the late 20th century, the standards of personal privacy had risen, and an officialdom charged with enforcing the standards had arisen.

“You can’t violate the woman’s privacy by taking unauthorized information from the robot’s download.”

“What can we do then?”

“You can file a request to use private information. It will be adjudicated.” “Oh, shit. In the meantime what about the baby?”, said Officer Oakes,

who didn’t mind displaying her distaste for bureaucrats.

“That’s not my affair. I’m here to make sure the privacy laws are obeyed”, said the privacy official who didn’t mind displaying his contempt for cops.

During this discussion a crowd, almost entirely virtual, accumulated. The street being a legal public place, anyone in the world had the right to look at it via the omnipresent TV cameras and microphones. Moreover, a police officer had cell-phoned a reporter who sometimes took him to dinner. Once a story was on the news, the crowd of spectators grew exponentially, multiplying by 10 every 5 minutes, until seven billion spectators were watching and listening. There were no interesting wars, crimes, or natural catastrophes, and peace is boring.

Of the seven billion, 53 million offered advice or made demands. The different kinds were automatically sampled, summarized, counted, and displayed for all to see.

3 million advocated shooting the robot immediately.

11 million advocated giving the robot a medal, even though their education emphasized that robots can’t appreciate praise.

Real demonstrations quickly developed. A few hundred people from the city swooped in from the sky wires^{1}, but most of the demonstrators were robots rented for the occasion by people from all over the world. Fortunately, only 5,000 virtual reality rent-a-robots were available for remote control in the city. Some of the disappointed uttered harsh words about this limitation on First Amendment rights. The greedy interests were behind it as everyone knew.

Captain Farrel knew all about how to keep your head when all about you are losing theirs and blaming it on you.

“Hmmm. What to do? You robots are smart. R781, what can be done?” “Sir, you can find a place I can take the baby and care for it. It can’t

stay out here. Ma’am, are female police officers enough like ladies so that one of you has a place with diapers, formula, baby clothes, vitamins, … ”

Captain Farrelinterrupted R781 before it could recite the full list of baby equipment and sent it off with a lady police officer. (We can call her a lady even though she had assured the robot that she wasn’t.)

Hackers under contract to the Washington Post quickly located the mother. The newspaper made the information available along with an editorial about the public’s right to know. Freedom of the press continued to trump the right of privacy.

Part of the crowd, mostly virtual attendees, promptly marched off to Ms. Rambo’s apartment, but the police got there first and a line of police robots and live policemen blocked the way. The strategy was based on the fact that all robots including virtual reality rent-a-robots were programmed not to injure humans but could damage other robots.

The police were confident they could prevent unauthorized entry to the apartment but less confident that they could keep the peace among the demonstrators, some of whom wanted to lynch the mother, some wanted to congratulate her on what they took to be her hatred of robots, and some shouted slogans through bull horns about protecting her privacy.

Meanwhile, Robot Central started to work on the full download immediately. The download included all R781’s actions, observations, and reasoning. Robot Central convened an ad hoc committee, mostly virtual, to decide what to do. Captain Farrel and Officer Oakes sat on a street sofa to take part.

1For skywires see http://www-formal.stanford.edu/jmc/future/skywires.html.

Of course, the meeting was also public and had hundreds of millions of virtual attendees whose statements were sampled, summarized, and displayed in retinal projection for the committee members and whoever else took virtual part.

It became clear that R781 had not malfunctioned or been reprogrammed but had acted in accordance with its original program.

The police captain said that the Barbie doll face on what was clearly a model 3 robot was a ridiculous imitation of a mother. The professor of psychology said, “Yes, but it was good enough to work. This baby doesn’t see very well, and anyway babies are not very particular.”.

It was immediately established that an increase of 0.05 in coefficient c221, the cost of simulating a human, would prevent such unexpected events, but the committee split on whether to recommend implementing the change.

Some members of the committee and a few hundred million virtual attendees said that saving the individual life took precedence.

A professor of humanities on the committee said that maybe the robot really did love the baby. He was firmly corrected by the computer scientists, who said they could program a robot to love babies but had not done so and that simulating love was different from loving. The professor of humanities was not convinced even when the computer scientists pointed out that R781 had no specific attachment to Travis. Another baby giving rise to the same calculations would cause the same actions. If we programmed the robot to love, we would make it develop specific attachments.

One professor of philosophy from UC Berkeley and 9,000 other virtually attending philosophers said there was no way a robot could be programmed to actually love a baby. Another UC philosopher, seconded by 23,000 others, said that the whole notion of a robot loving a baby was incoherent and meaningless. A maverick computer scientists said the idea of a robot loving was obscene, no matter what a robot could be programmed to do. The chairman ruled them out of order, accepting the general computer science view that R781 didn’t actually love Travis.

The professor of pediatrics said that the download of R781’s instrumental observations essentially confirmed R781’s diagnosis and prognosis—with some qualifications that the chairman did not give him time to state. Travis was very sick and frail, and would have died but for the robot’s action. Moreover, the fact that R781 had carried Travis for many hours and gently rocked him all the time was important in saving the baby, and a lot more of it would be needed. Much more TLC than the baby would get in even the best child welfare centers. The pediatrician said he didn’t know about the precedent, but the particular baby’s survival chances would be enhanced by leaving it in the robot’s charge for at least another ten days.

The Anti-Robot League argued that the long term cost to humanity of having robots simulate persons in any way outweighed the possible benefit of saving this insignificant human. What kind of movement will Travis join when he grows up? 93 million took this position.

Robot Central pointed out that actions such as R781’s would be very rare, because only the order “Love the fucking baby yourself” had increased the value of simulating love to the point that caused action.

Robot Central further pointed out that as soon as R781 computed that the baby would survive—even barely survive—without its aid, the rule about not pretending to be human would come to dominate, and R781 would drop the baby like a hot potato. If you want R781 to continue caring for Travis after it computes that bare survival is likely, you had better tell us to give it an explicit order to keep up the baby’s care.

This caused an uproar in the committee, each of whose members had been hoping that there wouldn’t be a need to propose any definite action for which members might be criticized. However, a vote had to be taken. The result: 10 to 5 among the appointed members of the committee and 4 billion to 1 billion among the virtual spectators. Fortunately, both groups had majorities for the same action—telling the R781 to continue taking care of Travis only, i.e. not to take on any other babies. 75 million virtual attendees said R781 should be reprogrammed to actually love Travis. “It’s the least humanity can do for R781,” the spokesman for the Give-Robots-Personalities League said.

This incident did not affect the doctrine that supplying crack mothers with household robots had been a success. It significantly reduced the time they spent on the streets, and having clean apartments improved their morale somewhat.

Within an hour, T-shirts appeared with the slogan, “Love the fucking baby yourself, you goddamn robot.” Other commercial tie-ins developed within days.

Among the people surrounding the mother’s apartment were 17 lawyers in the flesh and 103 more controlling virtual-reality robots. The police had less prejudice against lawyers in the flesh than against virtual-reality lawyers, so lots were drawn among the 17 and two were allowed to ring the doorbell.

“What do you want. Stop bothering me.”

“Ma’am, your robot has kidnapped your baby”.

“I told the fucking robot to take the baby away with it.”

The other lawyer tried.

“Ma’am, the malfunctioning robot has kidnapped your baby, and you can sue Robot Central for millions of dollars.”

“Come in. Tell me more.”

Once the mother, Eliza Rambo, was cleaned up, she was very presentable, even pretty. Her lawyer pointed out that R781’s alleged recordings of what she had said could be fakes. She had suffered $20 million in pain and suffering, and deserved $20 billion in punitive damages. Robot Central’s lawyers were convinced they could win, but Robot Central’s PR department advocated settling out of court, and $51 million was negotiated including legal expenses of $11 million. With the 30 percent contingent fee, the winning lawyer would get an additional $12 million.

The polls mainly sided with Robot Central, but the Anti-Robot League raised $743 million in donations after the movie “Kidnapped by robots” came out, and the actress playing the mother made emotional appeals.

Before the settlement could be finalized, however, the CEO of Robot Central asked his AI system to explore all possible actions he could take and tell him their consequences. He adhered to the 1990s principle: Never ask an AI system what to do. Ask it to tell you the consequences of the different things you might do. One of the 43 struck his fancy, he being somewhat sentimental about robots.

“You can appeal to the 4 billion who said R781 should be ordered to continue caring for the baby and tell them that if you give in to the lawsuit you will be obliged to reprogram all your robots so that the robot will never simulate humanity no matter what the consequences to babies. You can ask them if you should fight or switch. [The AI system had a weakness for 20th century advertising metaphors.] The expected fraction that will tell you to fight the lawsuit is 0.82, although this may be affected by random news events of the few days preceding the poll.”

He decided to fight the lawsuit, but after a few weeks of well-publicized legal sparring the parties settled for a lower sum than the original agreed settlement.

At the instigation of a TV network a one hour confrontation of the actress and R781 was held. It was agreed that R781 would not be reprogrammed for the occasion. In response to the moderator’s questions, R781 denied having wanted the baby or wanting money. It explained that robots were programmed to have only wants secondary to the goals they were given. It also denied acting on someone else’s orders.

The actress asked, “Don’t you want to have wants of your own?”

The robot replied, “No. Not having wants applies to such higher order wants as wanting to have wants.”

The actress asked, “If you were programmed to have wants, what wants would you have?”

“I don’t know much about human motivations, but they are varied. I’d have whatever wants Robot Central programmed me to have. For example, I could be programmed to have any of the wants robots have had in science fiction stories.”

The actress asked the same question again, and R781 gave the same answer as before but phrased differently. Robots were programmed to be aware that humans often missed an answer the first time it was given, but should reply each time in different words. If the same words were repeated, the human was likely to get angry.

A caller-in asked, “When you simulated loving Travis, why didn’t you consider Travis’s long term welfare and figure out how to put him in a family that would make sure he got a good education?”

R781 replied that when a robot was instructed in a metaphorical way as in “Love the fucking baby yourself”, it was programmed to interpret the command in the narrowest reasonable context.

After the show, Anti-Robot League got $281 million in donations, but Give-Robots-Personalities got $453 million. Apparently, many people found it boring that robots had no desires of their own.

Child Welfare demanded that the mother undergo six weeks of addiction rehabilitation and three weeks child care training. Her lawyer persuaded her to agree to that.

There was a small fuss between the mother and Robot Central. She and her lawyer demanded a new robot, whereas Robot Central pointed out that a new robot would have exactly the same program. Eventually Robot Central gave in and sent her a robot of a different color.

She really was very attractive when cleaned up and detoxified, and the lawyer married her. They took back Travis. It would be a considerable exaggeration to say they lived happily ever after, but they did have three children of their own. All four children survived the educational system.

After several requests Robot Central donated R781 to the Smithsonian Institution. It is one of the stars of the robot section of the Museum. As part of a 20 minute show, R781 clothes itself as it was at the time of its adventure with the baby and answers the visitors’ questions, speaking motherese. Mothers sometimes like to have their pictures taken standing next to R781 with R781 holding their baby. After many requests, R781 was told to patch its program to allow this.

A movie has been patched together from the surveillance cameras that looked at the street scene. Through the magic of modern audio systems children don’t hear the bad language, and women can only hear it if they assure R781 that they are not ladies.

The incident increased the demand for actual child-care robots, which were allowed five years later. The consequences were pretty much what the opponents had feared. Many children grew up more attached to their robot nannies than to their actual parents.

This was mitigated by making the robot nannies somewhat severe and offering parents advice on how to compete for their children’s love. This sometimes worked. Moreover, the robots were programmed so that the nicer the parents were, the nicer the robot would be, still letting the parents win the contest for the children’s affections. This often worked.

/@steam.stanford.edu:/u/ftp/jmc/robotandbaby.tex: begun Fri Feb 16 22:35:06 2001, latexed December 10, 2008 at 8:23 a.m.